A flywheel of radius 14 in is rotating at the rate of 1000 rpm. How fast does a point on the rim travel in fps?
1
Expert's answer
2020-11-17T10:34:07-0500

The angular speed of the flywheel in radians per second is:

"w=\frac{1000 ⋅ 2\pi}{60}=104.7 rad/s"

The linear spped of a point on the rim of the flywheel is:

"v=w\u22c5r=104.7\u22c514=1465.3m\/s"

1 m=3.28 feet

v=1465.3 m/s=4806.3 feet/s